Proposals for IRAM Telescopes

The deadline for submission of observing proposals on IRAM telescopes, both the interferometer and the 30m, is

18 September 2008, 17:00 MEST (UT + 2 hours)

The scheduling period extends from 1 Dec 2008 – 31 May 2009. Proposals should be submitted through our web–based submission facility. Instructions can be found on our web page at URL:

http://www.iram.fr/GENERAL/

submission/submission.html

Detailed information on time estimates, special observing modes, technical information and references for both the IRAM interferometer and the IRAM 30m telescope can be found on the above mentioned web page. The submission facility will be opened about three weeks before the proposal deadline. Proposal form pages and the 30m time estimator are available now.

Please avoid last minute submissions when the network could be congested. As an insurance against network congestion or failure, we still accept, in well justified cases, proposals submitted by:

– fax to number: (+33) 476 42 54 69 or by

– ordinary mail addressed to:

IRAM Scientific Secretariat, 300, rue de la piscine, F-38406 St. Martin d’Hères, France

Proposals sent by e–mail are not accepted. Color plots will be printed/copied in grey scale. If color is considered essential for the understanding of a specific figure, a respective remark should be added in the figure caption. The color version may then be consulted in the electronic proposal by the referees.

Soon after the deadline the IRAM Scientific Secretariat sends an acknowledgement of receipt to the Principal Investigator of each proposal correctly received, together with the proposal registration number. Note that the web facility allows cancelation and modification of proposals before the deadline. The facility also allows to view the proposal in its final form as it appears after re–compilation at IRAM. We urge proposers to make use of this feature as we always receive a number of corrupted proposals (figures missing, blank pages, etc.).

Valid proposals contain the official cover page, up to two pages of text describing the scientific aims, and up to two more pages of figures, tables, and references. The normal proposals should not exceed these 5 pages. Except for the technical pages for the interferometer, longer proposals will be cut. The new Large Observing Programmes (see contribution elsewhere in this Newsletter) have up to 4 pages for the scientific justification, plus cover page and 2 pages for supporting material.

The proposal template proposal.tex and the L^{A}TEX style file proposal.sty may be obtained from the IRAM web pages^{1} at URL ../GENERAL/-

submission/proposal.html. In case of problems,

contact the secretary, Fabienne Schicke, (e–mail: berjaud@iram.fr). Please, make sure that your proposals use the current form pages.

In all cases, indicate on the proposal cover page whether your proposal is (or is not) a resubmission of a previously rejected proposal or a continuation of a previously accepted interferometer or 30m proposal. We request that the proposers describe very briefly in the introductory paragraph (automatically generated header “Proposal history: ”) why the proposal is being resubmitted (e.g. improved scientific justification) or is proposed to be continued (e.g. last observations suffered from bad weather).

Do not use characters smaller than 11pt. This could render your proposal illegible when copied or faxed. If we notice any formal problems before the deadline, we will make an effort to contact the principal investigator and solve the problem together.

Applications for short spacing observations have been simplified. If the need for complementary 30m observations is evident already at the time when the PdB interferometer proposal is prepared, just note this need on the interferometer proposal. A separate proposal for the 30m telescope is not required. The blank form for interferometer proposals contains a bullet, labelled “short spacings” which should then be checked. The interferometer style file will prompt for an additional paragraph in which the scientific need for the short spacings should be described. It is essential to give here all observational details, including size of map, sampling density and rms noise, spectral resolution, receiver configuration and time requested.

A mailing list has been set up for astronomers interested in being notified about the availability of a new Call for Proposals. A link to this mailing list is on the IRAM web page. The list presently contains all users of IRAM telescopes during the last 2 years.

J.M. Winters & C. Thum

1 from here on we give only relative URL addresses. In the absolute address the leading two dots (..) should be replaced by the address of one of our mirror sites: http://www.iram.fr or http://www.iram.es.

Travel funds for European astronomers

IRAM participates in the RadioNet project, an initiative funded by the European Commission within the FP6 Programme to improve and encourage communication among astronomers of the European Community and associated countries. Transnational access (TNA) is the largest RadioNet programme and provides funding for travel expenses incurred by eligible users for carrying out their observations or reducing their data. As a partner of RadioNet, IRAM has now some limited TNA funds to pay travel expenses for eligible users. Detailed information about user eligibility, TNA contacts, policies and travel claims for the IRAM 30m telescope and Plateau de Bure Interferometer can be found on the RadioNet home page at

http://www.radionet-eu.org.

As the FP6 Programme comes to an end this year and the terms of its successor FP7 have not yet been finalized, the TNA support which IRAM can provide is uncertain after December 2008. This caveat therefore concerns the bulk of the coming winter semester. Eligible users will be contacted directly when the new terms are available.

Observers requesting TNA support will be asked to provide the necessary personal and professional information to IRAM. Funding through RadioNet should be acknowledged in publications resulting from TNA supported observations.

R. Neri & C. Thum

Call for Observing Proposals on the 30m Telescope

Summary

Proposals for three types of receivers will be considered for the coming winter semester:

1. the observatory’s set of four dual polarization heterodyne receivers centered at wavelengths of 3, 2, 1.3, and 1.1 mm.

2. the 9 pixel dual-polarization heterodyne receiver array, HERA, operating at 1.3 mm wavelength

3. The MAMBO–2 bolometer array with 117 pixels operating at 1.2 mm; the smaller MAMBO– 1 array with 37 pixels is kept as a backup.

Emphasis will be put on observations at the shorter wavelengths, but 3mm proposals are also encouraged inasmuch as they are suited for medium or low quality weather backup onservations. About 2000 hours of observing time are expected to be available.

The main news relevant for the coming winter semester are described here. Details of proposal formalities, instrumentation, observing modes, and estimation of observing time are described on the IRAM web site.

What is new?

In addition to the normal observing proposals, IRAM invites applications for special Large Observing Programmes (see the contribution by P. Cox elsewhere in this Newsletter). On the 30m telescope, these Large Observing Proposals are restricted for the coming winter semester to the bolometer and HERA instruments. The proposal cover page provides a checkbox for identifying a Large Programme.

The next generation single pixel heterodyne receiver for Pico Veleta, EMIR (Eight MIxer Receiver), consisting of dual-polarization 4 GHz bandwidth mixers operating at 3, 2, 1.3, and 0.9mm, will provide a boost in sensitivity and observing capabilities, fully justifying its installation as soon as possible during the coming winter semester. Installation and commissioning will take about 4 weeks. During installation, observation with HERA or MAMBO can still precede during night time, since the Nasmyth cabin optics are not affected at this stage. In view of the uncertain time scale and our lack of experience with EMIR, we request 30m proposer to use the performance of the current receivers for their estimate of observing time. Proposals scheduled after the installation of EMIR may see their time allocation adjusted accordingly.

An effort was made with EMIR to keep as much as possible of the frequency range below 83 GHz. As the final outcome will not be known before the proposal deadline, we recommend to interested astronomers to consider applying for these low frequencies now.

Remote observing is available from the IRAM offices in Granada and Grenoble, and from the remote stations in Madrid and Bonn. A remote station in Paris may also become available soon.

The full Call for 30m Proposals is available on the IRAM web site at the URL indicated above.

Clemens Thum & Carsten Kramer